A child sitting 1.90 m from the center of a merry-go-round moves with a speed of 2.15 m/s .

Part A: Calculate the centripetal acceleration of the child.

Part B: Calculate the net horizontal force exerted on the child. (mass = 28.5 kg )
Answer #1

The centripetal acceleration is towards the center and is given by a =v^2/r

=(2.15)^2 / 1.90

= 2.433 m/s^2

(b) Net force in horizontal is centripetal force acting towards the center

F = ma

=28.5× 2.433

=69.2405 N